OpenStudy (anonymous):

This is an introductory question for calculus 1: if given the interval notation of [2,infinity) what would that be convert to a inequality

OpenStudy (zzr0ck3r):

so x is allowed to be 2, but of course will NEVER be infinity. So we get \(2\le x < \infty\)

OpenStudy (zzr0ck3r):

\([2,\infty)\) means all the numbers between 2 and infinity. and the [ means we are allowed to be 2.

OpenStudy (anonymous):

Yea i got that but graphically how would that be represented?

OpenStudy (zzr0ck3r):

If they gave you \((2,\infty)\) it would have been \(2<x<\infty\)

OpenStudy (zzr0ck3r):



OpenStudy (anonymous):

Even if the equality had been 2 <= x <infinity it would be graphed as shown above

OpenStudy (anonymous):

Oh ok never mind i see what you were referring to thank you